# On the $\Gamma$-limit of weighted fractional energies

### v1

@@ Abstract.
$[u]_{s,p,f}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y$
as $s\to 1^{-}$ for $u\in L^{p}(\Omega)$, where $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$. Assuming that $(f_{s})_{s\in(0,1)}\subset L^{\infty}(\mathbb{R}^{d};[0,\infty))$ and $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$ are such that $f_{s}\to f$ in $L^{\infty}(\mathbb{R}^{d})$ as $s\to 1^{-}$, we show that $(1-s)[u]_{s,p,f_{s}}$ $\Gamma$-converges to the Dirichlet $p$-energy weighted by $f^{2}$. In the case $p=2$, we also prove the convergence of the corresponding gradient flows.
@@
Gagliardo seminorms, $\Gamma$-convergence, fractional gradient flows
@@ 1.1. Framework
We let $d\in\mathbb{N}$, $s\in(0,1)$ and $p\in[1,\infty)$. Given a nonnegative weight $g\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, our aim is to study the $\Gamma$-convergence as $s\to 1^{-}$ of the non-homogeneous (or weighted by $g$) $s$-fractional $p$-seminorm
$[u]_{s,p,g}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{\|x-y\|^{d+sp}}\,g(x)\,g(y)\,\mathrm{d}x\,\mathrm{d}y$ (1.1)
@@
The convergence as $s\to 1^{-}$ of (1.1) in the case $g\equiv 1$ —for which we use the shorthand $[\,\cdot\,]_{s,p}^{p}=[\,\cdot\,]_{s,p,1}^{p}$— has been deeply studied in recent years, both in the pointwise and in the $\Gamma$-sense. Since the literature is very vast, here we limit ourselves to a non-comprehensive list of results which are closer to the spirit of the present work.
The pointwise limit of the seminorm $[\,\cdot\,]^{p}_{s,p}$ as $s\to 1^{-}$ is a notable instance of the celebrated Bourgain–Brezis–Mironescu (BBM, for short) formula [BBM01, Dav02], yielding that $(1-s)[\,\cdot\,]_{s,p}$ converges to the Dirichlet $p$-energy up to a multiplicative constant. After the seminal contributions [BBM01, Dav02], the BBM formula has been extensively studied in several directions, see [Pon04a, Pon04b, DDG24] for more general results and [LS11, LS14] for extensions to arbitrary domains. We also refer to [Lud14-S, Lud14-p] for anisotropic fractional energies and to [DDP24, GS25] for sharp conditions for the validity of the BBM formula.
The $\Gamma$-convergence of $(1-s)[\,\cdot\,]_{s,p}^{p}$ to the Dirichlet $p$-energy as $s\to 1^{-}$ has been established in [CdLKNP23, DDP24] for the case $p=2$ and in [GS25] for every $p\in[1,\infty)$. We also refer to [BBM01, Pon04a] for similar results on bounded open sets.
@@
Beyond the case $g\equiv 1$, the asymptotic behavior of (1.1) and of similarly-defined energies has been studied for some particular weights, see [CCLP23, DL21] for the Gaussian framework and [K24] for weights depending on negative powers of the distance from the boundary.
The aim of the present paper is to investigate the asymptotic behavior of the weighted seminorm (1.1) as $s\to 1^{-}$ as the weight $f$ is allowed to vary with respect to the parameter $s$. In particular, our work provides a suitable extension of the results obtained in [CdLKNP23] as $s\to 1^{-}$ to the weighted setting.
Our main result, theorem 1.1 below, deals with the $\Gamma$-convergence of the energy (1.1) with respect to a uniformly converging family of weights $(f_{n})_{n\in\mathbb{N}}$ in $L^{\infty}(\mathbb{R}^{d};[0,\infty))$, whose limit $f$ is in $\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. Precisely, we prove that the $\Gamma$-limit is given by
$u\mapsto\begin{cases}\displaystyle K_{d,p}\|\nabla u\|_{p,f^{2}}^{p}=K_{d,p}\int_{\Omega}f^{2}\,\|\nabla u\|^{p}\,\mathrm{d}x&\text{for}\ p\in(1,\infty),\\[12.91663pt] \displaystyle K_{d,1}\|Du\|_{1,f^{2}}=K_{d,1}\int_{\Omega}f^{2}\,\mathrm{d}|Du|&\text{for}\ p=1,\end{cases}$
@@
see [BSZ24]*Lem. 2.1. Here and below, given a measurable function $u\colon\Omega\to\mathbb{R}$ on an open set $\Omega\subset\mathbb{R}^{d}$, we define $\tilde{u}\colon\mathbb{R}^{d}\to\mathbb{R}$ be such that $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$.
@@
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p>1$ or $u\in BV_{0}(\Omega)$ if $p=1$.
@@
    ($\Gamma$-liminf inequality) If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in L^{p}(\Omega)$, then
@@
    ($\Gamma$-limsup inequality) If $u\in L^{p}(\Omega)$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u^{n}\to u$ in $L^{p}(\Omega)$ and
@@
It is worth mentioning that the Lipschitz regularity assumption on the boundary of the reference set $\Omega$ can be dropped, up to changing the definition of the spaces $W^{1,p}_{0}(\Omega)$ for $p>1$ and $BV_{0}(\Omega)$ for $p=1$ in a suitable way. Indeed, instead of closing the space $C^{\infty}_{c}(\Omega)$ with respect to the $W^{1,p}$ seminorm, for our scopes it would be enough to define
[…]
for $p>1$ and, similarly,
[…]
for $p=1$. However, we do not insist on this technicality here and work with the usual definition of the spaces, since such generality is beyond the purpose of the present note.
@@ 1.2. Convergence of flows
In the case $p=2$, the $\Gamma$-convergence result obtained in theorem 1.1 can be complemented with a stability result for the corresponding parabolic flows associated to the energies, see theorem 1.2 below. Here and below, given a weight $g\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, we define the weighted Laplacian of $u\in H^{1}_{0}(\Omega)$ as
$(-\mathfrak{D})^{g}u=-2\operatorname{div}(g^{2}\nabla u)$ (1.6)
in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Moreover, given $u\in L^{2}(\Omega)$ such that $[u]_{s,2,g}<\infty$, we define the weighted fractional $s$-Laplacian of $u$ as
$(-\mathfrak{D})^{s,g}u(x)=4g(x)\,\mathrm{p.v.}\int_{\mathbb{R}^{d}}\frac{\tilde{u}(x)-\tilde{u}(y)}{\|x-y\|^{d+2s}}\,g(y)\,\mathrm{d}y$ (1.7)
again in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Note that, in the unweighted case $g\equiv 1$, up to a multiplicative constant, the operators (1.6) and (1.7) become the usual Laplacian and fractional $s$-Laplacian operators, respectively.
@@
Let $(f_{n})_{n\in\mathbb{N}}$, $f$, $\Omega$, and $(s_{n})_{n\in\mathbb{N}}$ be as in theorem 1.1. If $(u^{n}_{0})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ is such that $u^{n}_{0}\to u^{\infty}_{0}$ in $L^{2}(\Omega)$ for some function $u^{\infty}_{0}\in L^{2}(\Omega)$, $[\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}<\infty$ for every $n\in\mathbb{N}$, and
$\sup_{n\in\mathbb{N}}(1-s_{n})[\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}<\infty,$
@@ 1.3. Organization of the paper
The paper is organized as follows. The notation and some useful preliminary results are detailed in section 2. The proof of theorem 1.1 is given in section 3, while that of theorem 1.2 can be found in section 4.
@@ 2.1. Notation
The symbol $C(*,\cdots,*)$ indicates a generic positive constant that depends on ,,*$\cdots$,*∗ , ⋯ , ∗ only and may change from line to line.
@@
Let $p\in[1,\infty)$ and $g\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$. Given $m\in\mathbb{N}$ and $v\colon\Omega\to\mathbb{R}^{m}$, we let
$\|v\|_{p,g}=\left(\int_{\Omega}\|v(x)\|^{p}\,g(x)\,\mathrm{d}x\right)^{\frac{1}{p}}\in[0,\infty]$ (2.2)
@@
When $m=1$, we simply write $L^{p}_{g}(\Omega)$. We point out that if additionally $g$ takes values in $(0,\infty)$, under our standing assumptions on $g$ and $\Omega$, the spaces $L^{p}(\Omega)$ and $L^{p}_{g}(\Omega)$ are equivalent, with
$(\operatorname{ess\,inf}_{\Omega}g)\,\|v\|_{p}^{p}\leq\|v\|^{p}_{p,g}\leq\|g\|_{\infty}\|v\|^{p}_{p}.$ (2.3)
@@
$[u]_{s,p,g}=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,g(x)\,g(y)\,\mathrm{d}x\,\mathrm{d}y\right)^{\frac{1}{p}}$ (2.4)
for every $u\colon\Omega\to\mathbb{R}$ and we use the shorthand $[u]_{s,p}=[u]_{s,p,1}$.
Finally, we let $W^{1,p}_{0}(\Omega)$ for $p>1$, and $BV_{0}(\Omega)$ for $p=1$, be the closure of $C^{\infty}_{c}(\Omega)$ functions with respect to the Sobolev $p$-norm $u\mapsto\|u\|_{p}^{p}+\int_{\mathbb{R}^{d}}\|\nabla u\|_{p}^{p}\,\mathrm{d}x$. We also set
$\|Du\|_{1,g}=\int_{\mathbb{R}^{d}}g\,\mathrm{d}|Du|$ (2.5)
whenever $u\in BV_{0}(\Omega)$ and $g\in L^{\infty}(\mathbb{R}^{d};(0,\infty))$.
@@ 2.2. Compactness and characterization
We preliminary prove theorem 1.1 in the case $f_{n}=f$ for every $n\in\mathbb{N}$. We restate our result in this particular case for better clarity.
@@
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p\in(1,\infty)$ or $u\in BV_{0}(\Omega)$ if $p=1$.
@@
    ($\Gamma$-limsup inequality) If $u\in L^{p}(\Omega)$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u_{n}\to u$ in $L^{p}(\Omega)$ and
@@
The proof of the three statements (i), (ii), and (iii) of theorem 3.1 is split across sections 3.1, 3.2 and 3.3. The proof of theorem 1.1 is given in section 3.4.
3.1. Proof of theorem 3.1(i)
@@
Let $g\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{g}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{B_{\varrho}}\|\tau_{y}v-v\|_{L^{p}_{g}(E_{\|h\|})}^{p}\,\mathrm{d}y$ (3.4)
for every $v\in L^{p}(\mathbb{R}^{d})$, $h\in\mathbb{R}^{d}$, $\varrho\in(0,\|h\|]$ and every bounded open set $E\subset\mathbb{R}^{d}$, where $E_{\|h\|}$ is defined according to the notation in (2.1).
@@
$|V_{\varrho}(\xi)|^{p}\leq\frac{\omega_{d}}{\varrho^{d}}\,\|\varphi\|^{p}_{\infty}\int_{B_{\varrho}}|v(\xi)-\tau_{y}v(\xi)|^{p}\,\mathrm{d}y$ (3.7)
@@
Multiplying inequality (3.9) by $g(x)$ and integrating with respect to $x\in E$, the claim immediately follows by Fubini’s Theorem. We omit the simple details. ∎
@@
Let $g\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{g}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{B_{\|h\|}}\frac{\|\tau_{y}v-v\|_{L^{p}_{g}(E_{\|h\|})}^{p}}{\|y\|^{d+sp}}\,\mathrm{d}y$
@@
The proof of proposition 3.3 requires the following Hardy-type inequality, which is taken from [AdPM01]*Prop. 6.
@@
Actually, in the proof of proposition 3.3 we use the weaker estimate
@@
$\varphi_{v}(t)=\int_{\partial B_{t}}\|\tau_{y}v-v\|_{L^{p}_{g}(E_{\|h\|})}^{p}\,\mathrm{d}\mathcal{H}^{d-1}(y)$ (3.12)
@@
$\|\tau_{h}v-v\|_{L_{g}^{p}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t$ (3.13)
for some $C=C(d,p)>0$. We now multiply both sides of (3.13) by $\varrho^{-1+p-sp}$ and integrate in the interval $[0,\|h\|]$ with respect to $\varrho$, getting
$\|\tau_{h}v-v\|_{L_{g}^{p}(E)}^{p}\leq Cp(1-s)\frac{\|h\|^{p}}{\|h\|^{p-sp}}\int_{0}^{\|h\|}\frac{1}{\varrho^{d+sp+1}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t\,\mathrm{d}\varrho.$
@@
$\|\tau_{h}v-v\|_{L^{p}_{g}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{0}^{\|h\|}\frac{\varphi_{v}(t)}{t^{d+sp}}\,\mathrm{d}t$
@@
We are now ready to detail the proof of the compactness statement (i) in theorem 3.1.
Proof of theorem 3.1(i).
@@
By proposition 3.3 applied on $\Omega_{1}$ and by the previous inequality, we have
@@
$=C(1-s_{n})\|h\|^{s_{n}p}\int_{(\Omega_{1})_{\|h\|}}\int_{B_{\|h\|}(x)}\frac{|\tilde{u}^{n}(\xi)-\tilde{u}^{n}(x)|^{p}}{\|\xi-x\|^{d+s_{n}p}}f(\xi)f(x)\,\mathrm{d}\xi\,\mathrm{d}x$

### v2

@@ Abstract.
$u]_{s,p,f}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y,$
as $s\to 1^{-}$ for $u\in L^{p}(\Omega)$, where $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$. Assuming that $(f_{s})_{s\in(0,1)}\subset L^{\infty}(\mathbb{R}^{d};[0,\infty))$ and $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$ are such that $f_{s}\to f$ in $L^{\infty}(\mathbb{R}^{d})$ as $s\to 1^{-}$, we show that $(1-s)[u]_{s,p,f_{s}}^{p}$ $\Gamma$-converges to the Dirichlet $p$-energy weighted by $f^{2}$. In the case $p=2$, we also prove the convergence of the corresponding gradient flows.
@@
$\Gamma$-convergence, fractional gradient flows, Gagliardo seminorms, parabolic flows, weighted spaces
@@
This is a pre-print of an article published in Proc. Roy. Soc. Edinburgh Sect. A. The final authenticated version is available online at: https://doi.org/10.1017/prm.2025.10100
@@ 1.1. Framework
We let $d\in\mathbb{N}$, $s\in(0,1)$, and $p\in[1,\infty)$. Given a nonnegative weight $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, our aim is to study the $\Gamma$-convergence as $s\to 1^{-}$ of the non-homogeneous (or weighted by $f$) $s$-fractional $p$-seminorm
$u]_{s,p,f}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y,$ (1.1)
@@
The convergence as $s\to 1^{-}$ of (1.1) in the case $f\equiv 1$—for which we use the shorthand $\cdot\,]_{s,p}^{p}=[\,\cdot\,]_{s,p,1}^{p}$—has been deeply studied in recent years, both in the pointwise and in the $\Gamma$-sense. Since the literature is very vast, here we limit ourselves to a non-comprehensive list of results which are closer to the spirit of the present work.
The pointwise limit of the seminorm $\cdot\,]^{p}_{s,p}$ as $s\to 1^{-}$ is a notable instance of the celebrated Bourgain–Brezis–Mironescu (BBM, for short) formula [BBM01, Dav02], yielding that $(1-s)[\,\cdot\,]_{s,p}$ converges to the Dirichlet $p$-energy up to a multiplicative constant. After the seminal contributions [BBM01, Dav02], the BBM formula has been extensively studied in several directions, see [Pon04a, Pon04b, DDG24] for more general results and [LS11, LS14] for extensions to arbitrary domains. We also refer to [Fan25, Lud14-S, Lud14-p] for anisotropic fractional energies and to [DDP24, GS25] for sharp conditions for the validity of the BBM formula.
The $\Gamma$-convergence of $(1-s)[\,\cdot\,]_{s,p}^{p}$ to the Dirichlet $p$-energy as $s\to 1^{-}$ has been established in [BPS16] for every $p\in(1,\infty)$, in [CdLKNP23] only for $p=2$, and in [GS25] for every $p\in[1,\infty)$. We also refer to [BBM01, Pon04a] for similar results on bounded open sets.
@@
Beyond the case $f\equiv 1$, the asymptotic behavior of (1.1) and of similarly-defined energies has been studied for some particular weights, see [CCLP23, DL21] for the Gaussian framework and [K24] for weights depending on negative powers of the distance from the boundary.
The aim of the present paper is to investigate the asymptotic behavior of the weighted seminorm (1.1) as $s\to 1^{-}$ as the weight $f$ is also allowed to vary with respect to the parameter $s$. This is motivated by the recent interest in the extension of BBM-type formulas beyond the isotropic setting in order to address possible applications to non-isotropic frameworks [DDP24]. Besides, our $\Gamma$-convergence result can be interpreted as a suitable extension to the weighted setting of the ones obtained in [CdLKNP23, BPS16].
Our main result, Theorem˜1.1 below, deals with the $\Gamma$-convergence of the energy (1.1) with respect to a uniformly converging family of weights $(f_{n})_{n\in\mathbb{N}}$ in $L^{\infty}(\mathbb{R}^{d};[0,\infty))$, whose limit $f$ is in $\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. Precisely, we prove that the $\Gamma$-limit is given by
$u\mapsto\begin{cases}\displaystyle K_{d,p}\|\nabla u\|_{p,f^{2}}^{p}=K_{d,p}\int_{\Omega}f^{2}\,\|\nabla u\|^{p}\,\mathrm{d}x,&\text{for}\ p\in(1,\infty),\\[12.91663pt] \displaystyle K_{d,1}\|Du\|_{1,f^{2}}=K_{d,1}\int_{\Omega}f^{2}\,\mathrm{d}|Du|,&\text{for}\ p=1,\end{cases}$
@@
see [BSZ26]*Lem. 2.1. Here and below, given a measurable function $u\colon\Omega\to\mathbb{R}$ on an open set $\Omega\subset\mathbb{R}^{d}$, we define $\tilde{u}\colon\mathbb{R}^{d}\to\mathbb{R}$ be such that $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$.
@@
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p>1$ or $u\in BV(\Omega)$ if $p=1$.
@@
    ($\Gamma$-liminf inequality) If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then
@@
    ($\Gamma$-limsup inequality) If $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u^{n}\to u$ in $L^{p}(\Omega)$ and
@@ 1.2. Convergence of flows
In the case $p=2$, the $\Gamma$-convergence result obtained in Theorem˜1.1 can be complemented with a stability result for the corresponding parabolic flows associated to the energies, see Theorem˜1.2 below. Here and below, given a weight $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, we define the weighted Laplacian of $u\in H^{1}_{0}(\Omega)$ as
$(-\mathfrak{D})^{f}u=-2\operatorname{div}(f^{2}\nabla u)$ (1.6)
in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Moreover, given $u\in L^{2}(\Omega)$ such that $u]_{s,2,f}<\infty$, we define the weighted fractional $s$-Laplacian of $u$ as
$(-\mathfrak{D})^{s,f}u(x)=4f(x)\,\mathrm{p.v.}\int_{\mathbb{R}^{d}}\frac{\tilde{u}(x)-\tilde{u}(y)}{\|x-y\|^{d+2s}}\,f(y)\,\mathrm{d}y$ (1.7)
again in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Note that, in the unweighted case $f\equiv 1$, up to a multiplicative constant, the operators (1.6) and (1.7) become the usual Laplacian and fractional $s$-Laplacian operators, respectively.
@@
Let $(f_{n})_{n\in\mathbb{N}}$, $f$, $\Omega$, and $(s_{n})_{n\in\mathbb{N}}$ be as in Theorem˜1.1. If $(u^{n}_{0})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ is such that $u^{n}_{0}\to u^{\infty}_{0}$ in $L^{2}(\Omega)$ for some function $u^{\infty}_{0}\in L^{2}(\Omega)$, $\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}<\infty$ for every $n\in\mathbb{N}$, and
$\sup_{n\in\mathbb{N}}(1-s_{n})[\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}^{2}<\infty,$
@@ 1.3. Organization of the paper
The paper is organized as follows. The notation and some useful preliminary results are detailed in Section˜2. The proof of Theorem˜1.1 is given in Section˜3, while that of Theorem˜1.2 can be found in Section˜4.
@@ 2.1. Notation
The symbol $C(*,\cdots,*)$ indicates a generic positive constant that depends on $\cdots,*$ only and may change from line to line.
@@
Let $p\in[1,\infty)$ and $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$. Given $m\in\mathbb{N}$ and $v\colon\Omega\to\mathbb{R}^{m}$, we let
$\|v\|_{p,f}=\left(\int_{\Omega}\|v(x)\|^{p}\,f(x)\,\mathrm{d}x\right)^{\frac{1}{p}}\in[0,\infty],$ (2.2)
@@
When $m=1$, we simply write $L^{p}_{f}(\Omega)$. We point out that if additionally $f$ takes values in $(0,\infty)$, under our standing assumptions on $f$ and $\Omega$, the spaces $L^{p}(\Omega)$ and $L^{p}_{f}(\Omega)$ are equivalent, with
$(\operatorname{ess\,inf}_{\Omega}f)\,\|v\|_{p}^{p}\leq\|v\|^{p}_{p,f}\leq\|f\|_{\infty}\|v\|^{p}_{p}.$ (2.3)
@@
$u]_{s,p,f}=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y\right)^{\frac{1}{p}}$ (2.4)
for every $u\colon\Omega\to\mathbb{R}$ and we use the shorthand $u]_{s,p}=[u]_{s,p,1}$.
Finally, we let $W^{1,p}_{0}(\Omega)$ for $p>1$, be the closure of $C^{\infty}_{c}(\Omega)$ functions with respect to the Sobolev $p$-norm $u\mapsto\|u\|_{p}^{p}+\int_{\mathbb{R}^{d}}\|\nabla u\|_{p}^{p}\,\mathrm{d}x$, while $BV(\Omega)$ the weak closure of $C^{\infty}_{c}(\Omega)$ with respect to the Sobolev $1$-norm. We also set
$\|Du\|_{1,f}=\int_{\mathbb{R}^{d}}f\,\mathrm{d}|Du|,$ (2.5)
whenever $u\in BV(\Omega)$ and $f\in L^{\infty}(\mathbb{R}^{d};(0,\infty))$.
@@ 2.2. Compactness and characterization
We preliminarily prove Theorem˜1.1 in the case $f_{n}=f$ for every $n\in\mathbb{N}$. We restate our result in this particular case for better clarity.
@@
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p\in(1,\infty)$ or $u\in BV(\Omega)$ if $p=1$.
@@
    ($\Gamma$-limsup inequality) If $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u_{n}\to u$ in $L^{p}(\Omega)$ and
@@
The proof of the three statements (i), (ii), and (iii) of Theorem˜3.1 is split across Sections˜3.1, 3.2 and 3.3. The proof of Theorem˜1.1 is given in Section˜3.4.
3.1. Proof of Theorem˜3.1(i)
@@
Let $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{B_{\varrho}}\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}\,\mathrm{d}y$ (3.4)
for every $v\in L^{p}(\mathbb{R}^{d})$, $h\in\mathbb{R}^{d}$, $\varrho\in(0,\|h\|]$, and every bounded open set $E\subset\mathbb{R}^{d}$, where $E_{\|h\|}$ is defined according to the notation in (2.1).
@@
$|V_{\varrho}(\xi)|^{p}\leq\frac{\omega_{d}}{\varrho^{d}}\,\|\varphi\|^{p}_{\infty}\int_{B_{\varrho}}|v(\xi)-\tau_{y}v(\xi)|^{p}\,\mathrm{d}y,$ (3.7)
@@
Multiplying inequality (3.9) by $f(x)$ and integrating with respect to $x\in E$, the claim immediately follows by Fubini’s Theorem. We omit the simple details. ∎
@@
Let $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{B_{\|h\|}}\frac{\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}}{\|y\|^{d+sp}}\,\mathrm{d}y$
@@
The proof of Proposition˜3.3 requires the following Hardy-type inequality, which is taken from [AdPM01]*Prop. 6.
@@
Actually, in the proof of Proposition˜3.3 we use the weaker estimate
@@
$\varphi_{v}(t)=\int_{\partial B_{t}}\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}\,\mathrm{d}\mathcal{H}^{d-1}(y),$ (3.12)
@@
$\|\tau_{h}v-v\|_{L_{f}^{p}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t,$ (3.13)
for some $C=C(d,p)>0$. We now multiply both sides of (3.13) by $\varrho^{-1+p-sp}$ and integrate in the interval $0,\|h\|]$ with respect to $\varrho$, getting
$\|\tau_{h}v-v\|_{L_{f}^{p}(E)}^{p}\leq Cp(1-s)\frac{\|h\|^{p}}{\|h\|^{p-sp}}\int_{0}^{\|h\|}\frac{1}{\varrho^{d+sp+1}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t\,\mathrm{d}\varrho.$
@@
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{0}^{\|h\|}\frac{\varphi_{v}(t)}{t^{d+sp}}\,\mathrm{d}t,$
@@
We are now ready to detail the proof of the compactness statement (i) in Theorem˜3.1.
Proof of Theorem˜3.1(i).
@@
By Proposition˜3.3 applied on $\Omega_{1}$ and by the previous inequality, we have
@@
$\leq C(1-s_{n})\|h\|^{s_{n}p}\int_{(\Omega_{1})_{\|h\|}}\int_{B_{\|h\|}(x)}\frac{|\tilde{u}^{n}(\xi)-\tilde{u}^{n}(x)|^{p}}{\|\xi-x\|^{d+s_{n}p}}f(\xi)f(x)\,\mathrm{d}\xi\,\mathrm{d}x$